OpenStudy (anonymous):

Which of the following shows the cube root of 8? A. -2 B. 2 <--- C. -2 and 2 D. 4 @Dark_Force @yoyo123 @jtryon @jim_thompson5910 @Hero @TacoGod

OpenStudy (yoyo123):

Yup, 2*2*2 = 8

sammixboo (sammixboo):

8*2 doesn't equal 8 o.o

OpenStudy (anonymous):

^

sammixboo (sammixboo):

But yes... 2 cubed is 8

OpenStudy (anonymous):

You are correct. \[2^3=2\times2\times2=8\]

OpenStudy (anonymous):

Cubed is 'to the power of 3'. Square is 'to the power of 2.'

OpenStudy (anonymous):

And a negative won't work because:\[-2^3=-2\times-2\times-2=4\times-2=-8\]

OpenStudy (anonymous):

If x^2 = 196, what is the value of x? A. 4 B. 7 C. 14 <--- D. 392

OpenStudy (the_fizicx99):

It's asking to take the cube \(\ \sf root \). If you had a larger number it would be difficult to find your answer, therefore you have to use \(\ \sf \Large \sqrt[3]{8}\) which translates to \(\ \sf \Large 8^\frac{1}{3}\)

OpenStudy (anonymous):

You are correct again. \[14^2=14\times14=196\]

OpenStudy (the_fizicx99):

You want to find "x", so you have to get rid of that pesky square, so start by taking the inverse of a square, which is a square root. \(\ \sf \sqrt{196} \) = 14.  What you do to one side you do to the other to keep the equation balanced.